The length of a rectangular field is 50 feet less than its width, if the perimeter of the field is 840 feet, find the length of the field show your work
Answer by TimothyLamb(4379)

---
x = y - 50
2x + 2y = 840
---
put the system of linear equations into standard form
---
x - y = -50
2x + 2y = 840
---
copy and paste the above standard form linear equations in to this (matrix method) solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer1:
x= length = 185 ft
y= width = 235 ft
---
---
now solve by substitution:
---
2x + 2y = 840
2(y - 50) + 2y = 840
2y - 100 + 2y = 840
4y = 940
y = 235
x = 235 - 50
x = 185
---
answer2:
x= length = 185 ft
y= width = 235 ft
NOTE: matrix and substitution methods yield the same result
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php